RC meeting and Run Plan, Thursday November 10th
From SBS wiki
Return to the RC Meeting Calendar
RC Meeting Information
Time:at 3:30pm today
- Location: Counting house conference room
- Join ZoomGov Meeting
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
Accelerator and Hall Status
- After being down for more than 24 hours we finally got beam around 6pm last night. MCC sent us beam exclusively for an hour in the beginning. Simona started beam checkout.
- As soon as Hall C was up with 70uA, beam became very trippy. Around 10pm MCC took away beam for 20 mins to correct pathlength. [ https://logbooks.jlab.org/entry/4079522 ]
- During owl beam was relatively stable. Hall C didn't go above 60uA overnight.
- Simona finished C-hole alignment during owl. Currently she is doing IC calibration on C-optics. After that she will move to reference cell.
- The next pass change was scheduled to take place on 11/22 (two weeks from now). Mark has talked to the spokespersons and confirmed that we will stay at 3pass until December 19th. I talked to the PD afterwards and requested her to update the schedule.
- MCC is thinking about doing beam maintenance either on Tuesday or Wednesday.
- Frequent laser trips during swing and owl - Solved -- Bill
- One of the laser fiber temperature is higher than usual. Is it safe to run like that? - Target group decided to discuss offline and keep me posted
- Heater RTD signals are inconsistent. Currently running at lower temperature. What's the status? - Arun said the temperature difference between the top and bottom RTDs looking reasonable now.
- Calibration constant at the current kinematic setting. - Hunter will do EPR measurement to get the calibration constant as soon as beam checkout is done.
- Current Polarization: 10.2mV -> 12.2mV-> (BB ON + SBS OFF) |
- If holding field is changed for any reason, do NOT perform an NMR measurement.
- Hourly screenshots are necessary for the laser gui.
- Update from target team.
Any additional subsystem updates?
- Finish beam checkout - Simona [We are here. Should take ~2hrs or less if beam is stable]
- Get calibration constant for target polarization (EPR + NMR) - Hunter
- Do Ion chamber calibration on reference cell with H2 - Simona
- Take ~ 2 hours of data on H2 (147 psi, max current 45uA | ramp up in steps).
- Go to polarized He3 and start production.
- 45uA of beam current (ramp up slowly) | Use ps3=ps5=0, rest =-1 | One hour long runs
- Move the target to the pickup coil position and do an NMR sweep every 2/3 hours
- Return to step 1.
- Moller measurement - Sometime next week? According to Doug, Hall B is getting higher than expected polarization which may mean that we could get lower than expected. If that turns out to be the case then a Wien adjustment will be necessary.
- Optics runs with sieve in
- GEM luminosity study
Instructions for Shift Crew
- Keep an eye out for drifting BPM positions
- Log all DAQ crashes in the excel worksheet along with the reason for the crash
- If beam has been down due to accelerator issues perform a checkout of the beam profile by doing harp scans and check beam centering on target using the Carbon hole target, see Simona's procedure: procedure, If in doubt about whether you should go through the beam checkout procedure, call this number: 7578693131.
- Cross check the rate with what is expected: log
- Cross check with the expected ion chamber readings: log
Be sure to follow these procedures:
- Power down SBS, BB, and Detector HV before controlled access and allowing RadCon in the hall
- Communicate to the TO that the SBS and BB magnets will be powered down, so they can update the holding field
- See the beam recovery procedures at: HOW_TOs#Beam_Recovery_Procedure
- See the instructions to set the raster at: HOW_TOs#How_to_ask_MCC_to_set_the_desired_raster_size_of_6_mm
Experimental progress (for MCC planning meeting)
- 6.1% of accumulated charge for production data (13.55 Coulombs out of 232)
- 32% Experimental run time (46 out of 141 days)
- Long term "to do" items; not urgent, some can be done when 3He target not available
- Luminosity scans with GEMs. See procedure.
- Opportunistic BCM calibration (higher current) before pass change?
- INFN LV Scan and DAQ Stability - maybe 20-30 minutes if nothing else can run
- Work and testing on LV supplies was performed during opportunistic access. See log.
- Misc. GEM commissioning items: https://sbs.jlab.org/wiki/index.php/CommissioningPlan#BB.2FSBS_GEM_Commissioning_.28Continuation.29